### v1

# AgentHub: A Research Agenda for Agent Sharing Infrastructure

Abstract.
LLM-based agents are rapidly proliferating, yet the infrastructure for discovering, evaluating, and governing them remains fragmented compared to mature ecosystems like software package registries (e.g., npm) and model hubs (e.g., Hugging Face). Recent research and engineering works have begun to consider the requisite infrastructure, but so far they focus narrowly—on distribution, naming, or protocol negotiation. However, considering broader software engineering requirements would improve open-source distribution and ease reuse. We therefore propose AgentHub, a research agenda for agent sharing. By framing the key challenges of capability clarity, lifecycle transparency, interoperability, governance, security, and workflow integration, AgentHub charts a community-wide agenda for building reliable and scalable agent ecosystems. Our vision is a future where agents can be shared, trusted, and composed as seamlessly as today’s software libraries.
copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXconference: International Conference on Software Engineering; April 12–18, 2026; Rio De Janeiro, Brazilisbn: XXX-X-XXXX-XXXX-X/2025/11
1. Introduction
LLM-based agents are rapidly entering workflows, from scientific discovery (gottweis_towards_2025, ) to software engineering (zhang2024autocoderover, ). Unlike static software packages or pretrained models (JiangPTMReuse, ), Agents act with autonomy, compose tools dynamically, evolve (self-refine) over time, and can operate at scale (GoogleCloudAIAgents, ; He2025LLMMultiAgent, ). We believe that these attributes necessitate a new approach to sharing and composing the associated artifacts. As agent adoption grows, the lack of suitable infrastructure risks limiting both research progress and real-world impact.
In designing a registry for agent sharing, we can learn from the ecosystems for earlier kinds of software. Conventional software registries such as PyPI, npm, and Maven Central show the value of structured metadata, dependency graphs, and signed provenance (Zimmermann2019, ; schorlemmer2024signing, ). More recently, registries for pre-trained models, e.g., Hugging Face, expose artifacts and informal model cards, but in an effort to keep up with rapid change they omit normalized dependency and capability schemas, hampering reuse (JiangPTMReuse, ; jiang2024peatmoss, ; Jiang2025Naming, ). Meanwhile, emerging agent protocols, including the Model Context Protocol (MCP) (mcp-registry-repo, ; mcp-registry-blog, ), and the Agent Name Service (ANS) (ans, ), strengthen connectivity and naming but stop short of delivering a registry layer. The result: a fragmented landscape lacking features such as capabilities evidence and lifecycle status.
Figure 1. Conceptual view of AgentHub, illustrating how publishers, identities, and agent protocols might interact.
We envision AgentHub, a registry to support the production and consumption of software agents (Figure 1). An agent registry shares some requirements of earlier registries. However, unlike conventional artifacts, agents’ autonomy and dynamic composition demand transparent capability schemas and evidence (section 3.2.1); their continual evolution requires lifecycle visibility and fast revocation (section 3.2.2); cross-protocol operation calls for ecosystem interoperability (section 3.2.3); and large-scale, automated reuse raises openness, governance, and security concerns (sections 3.2.4 and 3.2.5).
Table 1. Aspects of existing software registries, and implications for AgentHub
Aspects Required Examples in Software Registries Implication For AgentHub
Metadata and dependency schema Package manifests (npm, PyPI, HF cards) encode machine-readable metadata and support versioned dependency graphs with auto-resolution (npm, Maven) (npm-trusted, ; pypi-pep458, ; hf-modelcards, ) Shared ontology (capabilities, I/O, protocols, provenance) with explicit agent–agent/service dependencies for reproducibility
Trust and provenance Signing and provenance (npm ECDSA; PyPI TUF) (npm-signatures, ; pypi-pep458, ) Signed manifests and reproducibility attestations
Governance and lifecycle management Open vs. curated submission models (npm, CRAN/app stores) (npm-2fa, ; cran-policies, ; cran-checklist, ) and update/revocation mechanisms (PyPI TUF) (pypi-pep458, ) Hybrid governance: open submissions with vetting for high-risk agents, plus explicit lifecycle states (active/deprecated/retired) and emergency removal paths
Quality signals Ratings, downloads, badges Stats, ratings, benchmarks, audit badges for selection
2. Background and Related Work
To support our vision, we analyze both pre-agent software registries and recent work on agent directory services.
2.1. Lessons from Pre-Agent Software Registries
An agent registry can draw lessons from earlier registries (Table 1).
Metadata: Programming language package registries such as npm, PyPI, Maven Central, and CRAN demonstrate the importance of structured metadata and explicit dependency graphs. Package manifests (e.g., package.json, POM.xml) encode versioning, licensing, and dependency constraints in machine-readable form, enabling automated resolution and reproducible builds (pypi-pep458, ; npm-trusted, ; hf-modelcards, ). SBOM standards require explicit declaration of components and dependency relationships for provenance and traceability (ntia-sbom-minimum, ; ntia-sbom-framing, ); emerging AI/ML BOMs extend the same idea to models, datasets, and configurations (xia2024trustsoftwaresupplychains, ; bennet2025SPDXAIBOM, ). Hugging Face relies primarily on model cards with limited dependency information (JiangPTMReuse, ). For example, while some cards reference required libraries, pretrained checkpoints, or paired datasets, these links are neither mandatory nor normalized into a dependency graph schema. This lack of standardized schemas leads to inconsistent naming and hampers automated reproducibility (Jiang2025Naming, ; JiangScored22, ; jiang2024peatmoss, ). The lesson for AgentHub is that agent metadata must go beyond identifiers to include standardized schemas that capture capabilities, input–output modalities, protocol requirements, and declared dependencies.
Provenance: Trust and provenance mechanisms are central to registry design. Maven Central requires every artifact to be signed with a PGP key; npm supports registry signatures and “trusted publishing” using OIDC (npm-trusted, ; npm-signatures, ); and PyPI is adopting The Update Framework (TUF) for signed metadata (pypi-pep458, ). For agents, provenance is especially critical because dynamic tool bindings and environment access amplify the risks of impersonation or poisoning. Accordingly, AgentHub should require signed metadata and reproducibility attestations for all entries.
Governance and Submission Policies: Governance models highlight trade-offs between openness and safety. Npm and PyPI accept broad participation with light pre-checks, while CRAN and app stores such as Apple’s App Store impose strict manual review  (apple-appstore-guidelines, ; npm-2fa, ). Ecosystems also implement revocation: app stores can remotely disable malicious apps, and PyPI can yank faulty releases (pypi-pep458, ). Notably, the leftpad incident in npm illustrated the ecosystem-wide disruption that can follow from a poorly governed removal (npm-leftpad, ; theregister-leftpad, ). Governance requires care and is non-obvious with autonomous agents.
Discovery and Quality Signals: Registries, model hubs, and extension marketplaces provide user ratings, download counts, verification badges, and metadata-rich model cards (hf-modelcards, ; hf-popularity, ). These signals allow users to identify reputable contributions at scale. AgentHub should adopt reputation systems such as usage statistics, audits, or benchmark results to complement technical metadata.
2.2. Related Work on Emerging Agentic Systems
Several prior works have targeted a related use case: using agents for the services they provide. However, these works have not considered the registry use case, where actors can go to identify agents and agent components. The Agent Name Service (ANS) proposes a DNS-style directory for agents, offering secure, protocol-agnostic naming and discovery (ans, ). The Agent Capability Negotiation and Binding Protocol (ACNBP) builds on ANS to enable secure capability negotiation among heterogeneous agents (acnbp, ). The NANDA Index introduces a decentralized, peer-to-peer index of agents with cryptographically verifiable “AgentFacts” attesting to capabilities and permissions (nanda, ). Similarly, the MCP Registry catalogs MCP servers and their tools (mcp-registry-repo, ; mcp-registry-blog, ), improving discoverability within the MCP ecosystem. Finally, curated marketplaces such as the ChatGPT Plugin Store or Alexa Skills Store show how policy-enforced ecosystems can scale with user trust (chatgpt-plugins, ; alexa-skill-ecosystem, ; alexa-skills-policy, ). Vendor SDKs are beginning to support an agent-make-agent pattern; for example, Anthropic’s Claude Agent can also generate orchestrated subagents (AnthropicAgentSDK, ; AnthropicSubagents, ).
In formulating AgentHub, we observe that the emerging set of capabilities provided by prior works are necessary but not sufficient for the registry use case and the agent-make-agent scenario. Addressing this demand requires new infrastructure, moving beyond technical protocols or metadata-only overlays to ensure transparency, interoperability, and accountability, which we outline subsequently in the research agenda (section 3).
3. Research Agenda
We distinguish challenges shared by all registries (section 3.1) from those unique to agents (section 3.2).
Figure 2. Research agenda for AgentHub, showing how six requirements (§3.2) encounter specific challenges, motivating research directions toward objectives of reproducibility, portability, resilience, and fair discovery.
3.1. Common Registry Challenges
As indicated in Table 1, lessons learned from mature registries set the baseline: entries must carry structured manifests and dependency graphs for reproducibility; publishers and artifacts must be authenticated with signed metadata and public transparency logs; namespaces and lifecycle actions (publish, deprecate, revoke) must be governed; and the system must remain open and usable through simple APIs. Proven machinery such as PURL-style identifiers and existing code and model hubs should be incorporated through integration and adaptation, not re-invention. Since nearly all mainstream registries are centralized and host both metadata and artifacts, we expect similar centralization to be valuable for agent-related artifact management. However, running agents requires substantial hardware, so directory services will likely be necessary for that use case (section 2.2).
Next we present our research agenda in Figure 2: What changes when a registry’s artifacts, and some of its actors, are Agents?
3.2. Agent-Specific Considerations
3.2.1. Capability Clarity and Evidence
Registries for conventional software rely on manifests to make artifacts understandable and reproducible, but agents need a richer contract. For autonomous agents that compose tools and other agents at runtime, manifests must express runtime permissions, preconditions, environment bindings, and protocol roles. A useful analogy is Android’s permission model (android_permissions, ), where apps declare capabilities in a manifest that the OS validates during installation and use. Similarly, AgentHub manifests could expose agent capabilities in machine-readable form, enabling tools to flag over-privileged or under-evidenced agents before adoption. Engineers and agents should be able to plan for behaviors, not just observe them.
Fragmentation and Duplication: As npm’s “micro-packages” created brittle dependency chains (kula2017impactmicropackagesempiricalstudy, ; Abdalkareem2017TrivialPackages, ), agent ecosystems may spawn “micro-agents” with overlapping functions. Because agents can autonomously query registries and compose others, duplication and fragility can spread dynamically, enlarging the attack surface and degrading reliability, especially at scale as agents evolve and reimplement overlapping capabilities.
Equivalence: Beyond functional duplication, a deeper challenge is determining when two agents are truly equivalent. At a syntactic level, the same agent may appear across multiple repositories, creating confusion about which copy is authoritative. At a semantic level, multiple agents may claim the same capabilities but diverge in behavior due to nondeterminism and evolution in models, changes in context, and even hardware selection. Addressing this requires more than metadata alignment: AgentHub should support persistent identifiers, cross-registry attestations, and re-executable evidence pipelines to assess whether two agents are truly equivalent.
Abuse of Ambiguity: Ambiguity and missing metadata enable attacks in software registries (JiangPTMReuse, ); agents face the same risk. Unclear manifests let adversaries mimic popular entries or collaborators, echoing typosquatting and account hijacks in npm (Zimmermann2019, ). Agents’ autonomous installation/generation can accelerate such propagation unless strong signing and evidence requirements are enforced.
Near-term priorities are (i) standardizing machine-readable capability schemas covering capabilities, modalities, protocol roles, and dependencies enforced at publish time (JiangPTMReuse, ; Jiang2025Naming, ); (ii) supporting lightweight, re-executable (idempotent) evidence pipelines linking claims to traces or benchmark runs across versions; and (iii) adding badges or similar checks as digital nudges (brown2019nudges, ).
Some further discussion from a security view is in section 3.2.5.
3.2.2. Lifecycle Transparency
Software registries expose version history and revocations; autonomous agents need richer lifecycle states–active, deprecated, rotated, retired, or revoked–with timestamps and rationales. Because agents evolve dynamically and can continue acting without human oversight, lifecycle visibility is essential for safe reuse and governance. Discovery should respect these states so outdated entries do not appear healthy by default, and mirrors should propagate state changes within bounded freshness windows.
Dependency Concentration and Abandonment: The leftpad incident showed how removing even trivial packages disrupted thousands of projects. Zimmermann et al. found that a small number of npm maintainer accounts control most packages (Zimmermann2019, ), while Zerouali et al. note that popularity often masks inactivity (Zerouali2019, ). Agents face parallel risks: if many depend on a single base agent or are generated by one entity, failure or compromise could cascade broadly. Abandonment may occur not only when humans leave but also when autonomous agents stop updating, leaving stale yet discoverable entries in circulation.
Future work should define lifecycle metadata standards with clear states, timestamps, and rationales, along with monitoring to detect abandonment or unexpected behavioral drift. Research must also address how deprecated or revoked agents should appear in discovery, who has authority to mark or revoke them, and how federated mirrors should coordinate state changes. Comparing agent lifecycles with traditional software lifecycles may reveal where new loops–such as agents participating in design and implementation–demand stronger provenance, transparency, and control.
3.2.3. Ecosystem Interoperability
Planners and orchestrators compose agents and tools across protocols; cross-protocol operation requires a shared metadata core with protocol-specific extensions so intent-based queries can compare agents without losing meaning. It also requires portable, signed manifests and SBOM-style dependency graphs so tools, models, datasets, and services remain traceable across registries (e.g., npm, PyPI, model hubs) and agent protocols. Many model hubs still rely on ad-hoc files (e.g., free-form config.json), hurting portability and automated reuse.
One Schema, Many Dialects: Descriptors in MCP (tools), A2A (roles/behaviors), and ACP (message schemas/policies) use different primitives and evolve at different speeds. Without a standardized, signed manifest, semantics may be lost in translation, caches drift across mirrors, and naive popularity metrics dominate cross-protocol rankings while ignoring evidence quality and freshness. Stable identifiers for referenced artifacts are also missing, making cross-registry links fragile.
Short-term priorities include a compact capability ontology and a canonical manifest with required fields (capabilities, I/O modalities, protocol bindings, permissions, SBOM-style dependencies) and optional per-ecosystem extensions. Declarative adapters can then map native descriptors to this core and back, validated via round-trip conformance tests. Stable cross-registry identifiers (e.g., Purl-style) should be introduced and tested end-to-end with npm, PyPI, and model hubs. Early benchmarks could measure cross-protocol discovery in terms of precision/recall for intent queries, preservation of evidence link, freshness under churn, and ranking fairness as ecosystems scale.
3.2.4. Openness and Governance
At agent scale, publishers include both humans and automated pipelines, so governance must keep namespaces open and verifiable while preventing automated spam, squatting, and opaque takedowns. Traditional software registries such as PyPI and npm succeeded not only by providing distribution infrastructure but by embracing openness: anyone could publish, namespaces were transparent, and governance processes were clear. For AgentHub, these properties are even more critical. Without mechanisms for open contribution, transparent review, and consistent namespace management, registries risk becoming closed silos controlled by a few vendors as ecosystems evolve.
Workforce and Concentration of Expertise: Many projects are maintained by a single person, creating bus-factor risk. Zimmermann et al. emphasize that the issue is less maintainer shortage than concentrated control (Zimmermann2019, ). Governance models differ: npm allows instant publication, while curated ecosystems like CRAN impose stricter checks.
Even if agents can take over some maintenance tasks, expertise may still concentrate in a few organizations or key models. This creates the risk that critical agents depend on too few people. To avoid such single points of failure, AgentHub governance must balance openness with safeguards: partial vetting for high-impact agents, incentives for broader participation, and clear processes to revoke unsafe agents.
Future work should investigate governance structures that combine openness with verifiability. Possible directions include decentralized namespace assignment rooted in ANS, community-driven policy boards for resolving disputes, and auditable logs of publication and revocation decisions. Comparative studies of centralized versus federated governance models can illuminate trade-offs in consistency, adoption, and resilience. Open questions include how to embed transparency in decision-making without sacrificing agility, and how to make governance mechanisms both fair across domains and enforceable at internet scale.
3.2.5. Trust and Security
Autonomous composition widens the attack surface: agents can install, call, or even generate other agents, so identity, provenance, and revocation must be machine-enforceable. Entries should use signed manifests, verified namespace control, and provenance for builds, models, and datasets. A useful precedent comes from the supply-chain domain: the SLSA v1.1 Verification Summary Attestation (VSA) (slsa-vsa, ) standardizes how to publish structured, signed evidence of checks performed on artifacts, and AgentHub could adopt an analogous mechanism to keep validation auditable and portable. Such evidence must be authenticated (with optional third-party attestations for sensitive cases), and revocation or key rotation must propagate quickly across mirrors.
Privacy and Vulnerability: In software registries, account hijacks and malicious updates already erode trust (Zimmermann2019, ). LLM-based agents add privacy-specific attack surfaces: misconfiguration or weak governance can leak sensitive data or intellectual property (wang-etal-2025-unveiling-privacy, ; zharmagambetov2025agentdamprivacyleakageevaluation, ); training-data exposure can reconstruct confidential content (carlini2021extractingtrainingdatalarge, ); and real incidents have leaked corporate data (samsung_chatgpt_forbes_2023, ). Multi-agent prompt injection can propagate malicious instructions and compromise collective decision-making (lee2024promptinfectionllmtollmprompt, ; shi2025promptinjectionattacktool, ). Privilege escalation, guardrail bypasses, weak output validation, and insufficient access control enable unauthorized tool use and code execution (narajala2025securinggenaimultiagentsystems, ). The attack frontier remains underexplored.
AgentHub can deploy defenses such as signed manifests, provenance attestations (pypi-pep458, ; npm-trusted, ), verified namespaces via ANS (ans, ), and protections against typosquatting and package-confusion attacks (Zimmermann2019, ; jiang2025confuguardusingmetadatadetect, ). AI-specific risks include deserialization exploits and prompt-injection. It is crucial to test whether defenses from human-mediated registries hold for autonomous agent systems. New threats call for zero-trust privilege separation, runtime checks for I/O behavior, and privacy-preserving audit pipelines to verify evidence without leaks (he2025redteamingllmmultiagentsystems, ; narajala2025securinggenaimultiagentsystems, ; samsung_chatgpt_forbes_2023, ; wang-etal-2025-unveiling-privacy, ).
3.2.6. Discovery and Workflow Integration
In an agent ecosystem, the “users” of the registry are both developers and agents that act autonomously at scale, so discovery must be programmatic, rank by capability-and-evidence fit rather than popularity, and integrate directly into planning, CI/CD, and orchestration loops.
Keyword Search and Popularity Bias: Agents that recommend or install one another can create feedback loops that instill mediocre or unsafe entries. Experience from software registries highlights how naive signals mislead: Zerouali et al. show that different popularity metrics in npm yield inconsistent results (Zerouali2019, ), underlining how reliance on downloads or stars can mislead users; and Jones et al. observed that model popularity on HuggingFace correlates strongly with documentation quality (jones2024knowhuggingfacesystematic, ), suggesting that discovery in AgentHub would similarly benefit from incentives for high-quality documentation. Humans can partly correct for this by reading docs. Agents benefit greatly if those signals are captured as structured metadata and verifiable evidence. Without workflow integration, both humans and agents may fall back to ad-hoc search, undermining adoption and reproducibility.
AgentHub can improve current practice by combining keyword search with structured metadata and evidence linked to benchmarks, incorporating lifecycle states to avoid unsafe or outdated agents. Evaluation should identify which signals such as metadata and documentation quality ensure reliable discovery. Discovery should match user goals to agents by capabilities and evidence, ensuring that popularity does not overshadow quality. This resembles work on recommendation systems for software libraries like Code Librarian, which uses contextual signals to suggest packages (tao2023codelibrarian, ). Research benchmarking discovery is necessary to measure precision, recall, and resilience under ecosystem changes.
4. Future Plans
4.1. Roadmap and Evaluation
We suggest that the community prioritize first the design of AgentHub’s core mechanisms: standardized capability schemas and evidence pipelines, lifecycle metadata for transparency, canonical manifests for interoperability, and basic governance structures. We should seed AgentHub with a small, diverse set of agents and provide a lightweight CLI/SDK so discovery fits ordinary workflows. Evaluation can emphasize usability and reliability, while also exploring which metrics are most appropriate. Security efforts can focus first on known risks such as typosquatting, prompt injection, and account compromise.
Longer term, the community must tackle challenges such as detecting and mitigating emergent multi-agent attacks, ensuring dynamic provenance and lifecycle accountability, and integrating AgentHub with broader software and model registries. Governance may need to evolve toward federated models with auditable processes, while discovery might mature into recommendation systems that balance intent, evidence quality, and fairness.
To support continuous measurement and adaptation, the operators of AgentHub should provide infrastructure to collect user feedback and registry-level longitudinal data, enabling future studies of common problems and ecosystem evolution. As AI supply chain standardization advances (bennet2025SPDXAIBOM, ; owasp-aibom, ), AgentHub should remain aligned with emerging specifications such as AIBOM. In parallel, as the agentic ecosystem expands, the AgentHub operators will refine the interface and reinforce integration and access controls to ensure secure, seamless interaction with new agents.
4.2. Positioning AgentHub if AGI Centralizes Orchestration
Our analysis of AgentHub is grounded in the current landscape and capabilities of agents. No one knows the future, but we acknowledge that increasingly capable foundation models may internalize a great deal of the coordination and planning that motivates today’s agentic and multi-agent systems.
We view this not as an end to multi-agent infrastructure but as a shift in where coordination lives and, importantly, a shift that heightens the need for transparent capability schemas, lifecycle visibility, and governance (section 3.2). First, economics and operational constraints make a purely “AGI-first” strategy brittle. Highly general models are likely to remain expensive in compute, energy, and latency and they will not be uniformly deployable across every environment that needs automation. Many production workflows will continue to mix a capable planner with specialized, composable executors that are cheaper to run, easier to colocate with data, and better aligned with SLA and locality requirements. If AgentHub exposes cost, latency, and fidelity signals alongside capabilities and evidence, discovery can optimize for fitness under budget and policy, rather than only for raw model capacity.
Acknowledgments
Davis acknowledges support from NSF awards #2343596, #2537308, and #2452533. Thiruvathukal and Läufer acknowledge support from NSF award #2343595. Thiruvathukal acknowledges support from NSF award #2537309.